The area of some shape varies directly with the square of its side. If a shape with 6 cm sides has an area of 216 cm2, what would be the area of a shape of the same type with sides of 8 cm?

Answer by TimothyLamb(4379)

---
a = Cxx
C = a/xx
---
C = 216/(6*6)
C = 6
---
a = Cxx
a = 6*8*8
a = 384 sq.cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php